OpenStudy (anonymous):

Write the equation in slope-intercept form of the line that has a slope of 6 and contains the point (1, 1).

OpenStudy (anonymous):

\[y-y_1=m(x-x_1)\] \[y-1=6(x-1)\] \[y-1=6x-6\] \[y=6x-5\]

OpenStudy (anonymous):

that is slope intercept form. slope is 6, intercept is (0,-5)

OpenStudy (anonymous):

hello amistre

OpenStudy (amistre64):

yes, it is: \[6(x-x_o)+y_o\]

OpenStudy (amistre64):

howdy sat :)

OpenStudy (anonymous):

my new goal is to keep up with half of your medals. think i can do it?

OpenStudy (amistre64):

its gonna be tough, but you can try ;)

OpenStudy (anonymous):

\[\frac{\text{amistre}}{2}=\text{sat73}\]